OpenStudy (firejay5):

Solve each equation. Then check your solution. Show and explain so I can better understand it.

OpenStudy (firejay5):

\[\frac{ x }{ 2 } - \frac{ 1 }{ 3 } = \frac{ x }{ 3 } - \frac{ 1 }{ 2 }\]

OpenStudy (agent0smith):

Find the common denominator for all the fractions, then you can multiply both sides by that number.